OpenStudy (anonymous):

a rectangular container with height 24 in contains 44.9 gal of water when its full. a stick is placed diagonally from one corner of the upper.if BC= 10 in ,AC=25 in and b is a point indicating water level find the height and volume of the liquid in the container

OpenStudy (jacob902):

Please find the solution as below. i) Let AD be the height of the vessel = 24 in ii) Draw an altitude BE from B on to the base. iii) The triangles ACD & BCE are similar. So, BF/AD = BC/AC ==> BF/24 = 10/25 = 2/5 ==> BF = (2/5)*24 = 48/5 = 9.6 in So height of the water level = 9.6 inches. iv) Volume of water in the vessel = 44.9*9.6/24 = 17.96 gallons. If you are satisfied with the above solution kindly acknowledge, need not be BA - just with a simple word of Thanks. More than 90% of the persons who ask question are not even obliged to acknowledge with a simple thanks, which I believe keeps many from providing solutions. REQUEST: Those who wish to rate this solution with TD, you are welcome; but only one request, kindly give the reasons for the same in the comments column, which will make me to correct mistake and learn more. Thanks for your cooperation.

OpenStudy (mathstudent55):

@Jacob902 How can you possibly answer this question when the posted problem is incomplete and there is no figure to help understand the situation? What does this sentence mean? "a stick is placed diagonally from one corner of the upper." Also, how do you know where all those points are and which triangles are similar if we can't see any of them?

OpenStudy (jacob902):

i) Let AD be the height of the vessel = 24 in ii) Draw an altitude BE from B on to the base. iii) The triangles ACD & BCE are similar. So, BF/AD = BC/AC ==> BF/24 = 10/25 = 2/5 ==> BF = (2/5)*24 = 48/5 = 9.6 in So height of the water level = 9.6 inches. iv) Volume of water in the vessel = 44.9*9.6/24 = 17.96 gallons. If you are satisfied with the above solution kindly acknowledge, need not be BA - just with a simple word of Thanks. More than 90% of the persons who ask question are not even obliged to acknowledge with a simple thanks, which I believe keeps many from providing solutions. REQUEST: Those who wish to rate this solution with TD, you are welcome; but only one request, kindly give the reasons for the same in the comments column, which will make me to correct mistake and learn more. Thanks for your cooperation.

OpenStudy (anonymous):

thank you